OpenStudy (anonymous):

sec^2x dx/(secx+tanx)^2. integration from 0 to pi/2. Can someone please guide me to the answer

OpenStudy (mimi_x3):

Is it: \[\large \int\frac{sec^{2}xdx}{(secx+tanx)^{2}}\]?

OpenStudy (anonymous):

@Mimi_x3 Couldn't understand you. What do you want to know?

OpenStudy (mimi_x3):

Is that the integral? Just to make sure I don't want to do a integral which is not what you wrote.

OpenStudy (anonymous):

yeah. that is the integral. Lower limit is 0 and upper limit is pi/2.

OpenStudy (mimi_x3):

hm..would this work? \[\int\limits\frac{\sec^{2}xdx}{(\sin^{2}x+2sinx+1)\sec^{2}x} => \int\limits\frac{dx}{\sin^{2}x+2sinx+1} \]

OpenStudy (mimi_x3):

looks like a Weierstrass substitution http://en.wikipedia.org/wiki/Weierstrass_substitution

OpenStudy (anonymous):

then how to solve the equation in the denominator??

OpenStudy (anonymous):

Please solve this: